OpenStudy (anonymous):

Solve for L: A = L • W • H L = A times W over H L = W times H over A L = A over quantity W times H L = W over quantity H times A

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

@anonymous_user A little help here :)

OpenStudy (anonymous_user):

I think it's A. L = A times W over H

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

may you help me with one more @anonymous_user

OpenStudy (anonymous_user):

I can try. :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F. : C = 5 over 9(F – 32) Which of the following formulas correctly solves for F? F = 9 over 5C + 32 F = 9 over 5C – 32 F = 9C + 32 over 5 F = 9C - 32 over 5

OpenStudy (anonymous_user):

C, maybe.

OpenStudy (anonymous):

ok tthanks

OpenStudy (mathstudent55):

A = L • W • H To solve for L, you need to divide both sides by WH, so the answer is L = A/(WH)

OpenStudy (mathstudent55):

C = (5/9)(F – 32) First, divide both sides by 5/9 which is the same as multiplying both sides by 9/5 to get: (9/5)C = F - 32 Now add 32 to both sides: F = (9/5)C + 32